One example of a first order reaction is the isome
Answer #1

Using the first order kinetics equation

ln(Ao/At) = kt

ln(Ao/At) = 4.403 * 10^(-4) * 1.50 * 60 * 60

ln(Ao/At) = 2.37762

Taking antilog both sides we get

Ao/At = e^(2.37762)

At = 1/e^(2.37762) = 0.09277 grams